OpenStudy (anonymous):

Erica drove from Ottawa 80 km/h.Julie left after 1 hr and drove along the same road at 100km/h.How far from ottawa julie overtake.

OpenStudy (anonymous):

distance = velocity * time your looking for when did Erica and Julie drive the same distance so you have to even the time erica: d= 80*t Julie (began after one hour) d=100*(t-1) d=d there you go...